OpenStudy (anonymous):

Solve. x2 + 6x = 24 -3 ± sqrt 21 -3 ± sqrt 34 -3 ± sqrt 15 -3 ± sqrt 33

OpenStudy (anonymous):

\[x^2+6x=24\]-24 -24\[x^2+6x-24=0\]\[(-b+-\sqrt{b^2-4ac})/2a\]\[((-6)+-\sqrt{(6)^2-4(1)(-24)})/2(1)\]\[(-6+-\sqrt{36+96})/2\]\[(-6+-\sqrt{132})/2\]\[(-6+-(11.5))/2\]\[(-6+11.5)/2=2.75\]\[(-6-11.5)/2=-8.75\]

OpenStudy (anonymous):

Answer: \[-3\pm \sqrt{33}\]